OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

8x - 12 = 4x +24 8x - 4x = 24 +12 4x = 36 x = 9

OpenStudy (anonymous):

thank u